# Dynamics as Prompts: In-Context Learning for Sim-to-Real System Identifications

### v1

@@ I Introduction
In-context learning has gained traction as a method for adjusting model behavior without gradient updates, widely used in Natual Language Processing (NLP) [16] and recently applied in robotics to improve generalization [17, 18, 19]. For example, Xu et al. [19] enhanced the Decision Transformer (DT) [20] by using new task demonstrations as prompts for online adaptation. Most current in-context learning approaches focus on adapting policies when rewards or expert demonstrations changes, assuming fixed environment dynamics [18, 21]. While different dynamics could be framed as diverse tasks in a multi-task RL setting, more than it’s a counter-intuitive setting, it also becomes impractical with a high-dimensional continuous environment parameter space, requiring many tasks to capture the full range of behaviors. In this paper, we explore a novel question: “Can we adapt simulation environment parameters using the in-context learning paradigm?” Our goal is to eliminate the optimization loop in SysID, in order to accelerate the parameter estimation process by incorporating the in-context learning ability of transformer models.
We introduce in-Context AdaPTation modUle for sim-to-REal system identification, or CAPTURE, to bridge the sim-to-real gap. CAPTURE aims to dynamically adjust the environment parameters online to align simulated and real-world trajectories using next-token prediction based on past interaction data, which includes simulated trajectories, actions, environment parameters, and real-world trajectories as shown in Figure 1. CAPTURE frames the SysID problem as an in-context learning formulation, treating interaction histories as “context.” Unlike existing techniques [4, 22, 23] that rely on short state-action history, CAPTURE aim to learn the complex SysID search process itself through rich and multi-episodic interaction history data. Beyond learning single-step expert parameter matching behaviors, longer interaction histories enables the learned SysID causal transformer to capture a better dynamics representation of the environments. By incorporating in-context learning, CAPTURE provides a smoother and more accurate prediction of subsequent environmental parameters and dynamic behaviors.
@@ III Methodology
In this work, we take a different approach from traditional in-context policy adaptation. Instead of directly adapting the task policy, we focus on using the history of environment parameter updates to estimate the next-iteration environment parameters such that the simulation and real-world performance can be aligned. We assume that the sim-to-real performance gap will narrow as the discrepancy between the simulation and real-world environment parameters decreases. We start with a description of the problem formulation in Section III-A. Following with three key modules in our pipeline: Section III-B describes the task policy training, Section III-C describes how we generate efficient source-to-target adaptation iterations, and Section III-D defines different components in the SysID causal transformer structure.
@@ III-A Problem Formulation
In this section, we define the task of predicting accurate simulation environment parameters to align simulated dynamics with real-world environments. We outline how to model the interaction between dynamics behaviours and environment parameters as a sequence, forming the training data for SysID causal transformer models. We begin by introducing the simulation parameters, followed by the task policy and data generation notations, and conclude with the SysID causal transformer notations for domain adaptation.
Environment Parameter Space. We define the task-related environment parameter space $\epsilon\in\mathcal{E}$ that parameterized quantities such as the center of mass and sliding frictions. We also assume that the environment parameter space $\mathcal{E}$ is finite and bounded, encompassing properties of different objects. We modify the environment parameters with Robosuite [42], which provides API for modifying the environment parameters through Python code.
SysID Causal Transformer and Interaction Histories. During the SysID causal transformer and data collection setting, we use previous SysID iterations as context, including simulated state trajectories $\tau^{sim}=\{s^{sim}_{0},s^{sim}_{1},\dots,s^{sim}_{T}\}$, real state trajectories $\tau^{real}=\{s^{real}_{0},s^{real}_{1},\dots,s^{real}_{T}\}$, rollout action $a\sim\pi(a|s_{0},\epsilon)$, and the past environment parameters $\epsilon$. A robust SysID process explores complex parameter behaviours. We aim to capture this behavior using a causal transformer that infers parameters from past interactions. Following [17], we treat these sequential interactions as history, where current environment parameters depend on previous SysID iterations. Formally, we define the history as:
@@
where $h_{i}$ is the history containing the past $k$ iterations at $i$-th iteration. Our goal is to learn a causal transformer such that it can replicate the SysID process given history. We define the SysID causal transformer, $P_{\theta}$, with the objective of modeling the distribution of simulation parameters conditioned on the history. This encourages the simulated trajectories $\tau^{sim}$ eventually behave close to real-world trajectories, $\tau^{real}$, bridge the sim-to-real gap.
Given an ideal search strategy that successfully adapts to the target environment parameters, our goal is to learn the underlying search capabilities from this process by predicting the next iteration in the history. The optimization objective can be formalized as:
@@ III-B Environment-Conditioned RL Training
The environment-conditioned RL task policy $\pi(a\mid s_{0},\epsilon)$ is trained to adapt to varying environment parameters $\epsilon\in\mathcal{E}$. For each episode, $\epsilon$ is sampled uniformly from the parameter space $\mathcal{E}$. Within the episode, the agent selects an action $a$ from $\pi(a\mid s_{0},\epsilon)$, considering the initial state $s_{0}$ and environment parameter $\epsilon$. This action is executed, producing a state trajectories $\{s_{1},s_{2},\ldots,s_{T}\}$ and a reward $r$. Each episode $\{a,r,s_{0},\epsilon\}$, is stored in the replay buffer. After certain episodes, the policy is updated using Soft Actor-Critic (SAC) [43], refining actions for smoother domain adaptation with predicted parameters.
@@ III-C Source-to-Target SysID Iteration Generation
In the data generation process, we develop source-to-target adaptation transitions that mimic sim-to-real adaptation. Each iteration includes four elements: the current simulation parameter $\epsilon_{i}$, the policy action $a_{i}$, the simulated trajectories $\tau^{source}_{i}$, and the collected target environment trajectories $\tau^{target}_{i}$ under the same action $a_{i}$. The trajectories and actions are obtained through simulation rollouts using an environment-conditioned task policy.
In simulation, both source and target values are known, allowing for direct single-step mapping from source to target. However, this approach often performs poorly in real-world deployment when the target’s dynamics representation (state trajectories) lacks sufficient detail. Rather than learning a single-step mapping, we focus on learning a search algorithm that finds the target environment parameter with dynamic representations. The duration of the parameter iteration history $K$ indicates the number of iterations that we pre-defined to generate a complete transition sequence from $\epsilon^{source}$ to $\epsilon^{target}$. We pick a transition number $K=7$ during data generation.
@@
2:Choose parameter transition iteration length $K$
@@ III-D SysID Causal Transformer
Given the collected SysID parameter transition histories, $\mathcal{D}$, our goal is to distill the binary search process through parameter transition sequences with length $K$, where each iteration represents an adaptation iteration. The model predicts the next environment parameter $\hat{\epsilon}_{i+1}$ at iteration $i$ using a next-token prediction framework with a shifted input setup [45]. We sample a multi-episode window of size $k$ from $\mathcal{D}$, where $k$ is a subsequence of the full $K$ iterations. The SysID causal transformer processes this history to predict the next parameter. Each iteration block contains $2+2T$ tokens: one action, one parameter, and $T$ state trajectory tokens for both simulated and real rollouts.
@@ IV-A Experimental Setups
In table air hockey, we test the scalability of CAPTURE with a higher-dimensional parameter and action space [15, 48]. This task requires precise control and adaptability to match simulated and real-world dynamics. Tunable environment parameters are listed in Table I, with setups shown in Figure 4.
@@ IV-B Baselines and Ablations
Ablations on Different Data Generation Approaches. We modify the data generation module to demonstrate the effectiveness of our distilled searching algorithm over others, including linear interpolation (linterp) and the standard binary search method without randomness (binary), while selecting the next iteration parameters. Linear interpolation randomly selects $K$ points between source and target environment parameters and orderly constructs the SysID transition. The standard binary search method [49] follows a similar setting as ours. However, it does not consider the random beta distribution, it only selects the middle point between the upper and lower bound.
@@ IV-C Sim-to-Sim SysID Evaluation
Object Scooping Sim-to-Sim Evaluation. In the sim-to-sim transfer, we evaluated the normalized context differences, which are one-dimensional in this setting, as shown in Figure 5. Since we use an angle classifier for smoother real-world deployment, reporting trajectory differences would be meaningless, as the trajectory here is represented by a label. Instead, we measure the task’s success rate, defined as lift the object with label (0). Figure 5 shows that CAPTURE achieves a success rate 50% higher than other SysID methods and 70% higher than the DR approach. This is expected, as the baselines lack historical interaction data, making identification only dependent on current scooping points. In contrast, CAPTURE uses a rich previous interaction history, allowing it to gradually narrow down the center of mass search space.
@@
Table Air Hockey Sim-to-Sim Evaluation. In Figure 6, CAPTURE offers better parameter estimation with smoother and more accurate adaptation curves. In scenarios where environment parameters require rollout histories, baselines struggle due to their inability to account for historical interactions. For instance, while the ED method might successfully detect the left wall after hitting it, it tends to forget earlier right wall interactions. This short-term memory leads to faster adaptation in simple environments but falls short in more complex ones. In dynamic settings, where SysID needs to identify parameters on both sides for sustained task performance, maintaining a history of parameter updates becomes critical, as it informs subsequent iterations.
@@
Ablation results. In object scooping experiments, we observe that the linear interpolation approach converges more slowly in terms of adaptation iterations, as shown in Figure 5. Due to limited exploration, it hinders performance. Figure 6 shows that CAPTURE + linear interpolation follows a near-linear sim-to-real transition until iteration 7, closely matching the dataset’s transition history. However, it struggles to establish a robust search process due to overfitting to linear interpolated transition histories. Except for left damping parameter, no significant performance differences are seen between randomized binary search (ours) and standard binary search. With added randomized, it did not hinder the estimation performance, whereas it learned a more robust adaptation process. Figure 6 also illustrates that CAPTURE can smooth the adaptation process using history, regardless of SysID accuracy.
@@ V Conclusion
This paper introduces a novel in-context learning approach to bridge the sim-to-real gap in robotic tasks by adjusting environment parameters online. By leveraging interaction histories as context, we enable dynamics adaptation to real-world environments without requiring model updates. Evaluated in scooping and table air-hockey tasks, our method outperforms traditional approaches such as domain randomization and TuneNet, reducing the sim-to-real gap and improving both sim-to-sim and sim-to-real performance. The approach leverages historical multi-episode data to infer system parameters and provide a better real-world dynamics prediction. While our method demonstrates strong performance, it still requires to train a new SysID model for new task environments. Nonetheless, the framework offers a more efficient and accurate solution for real-world deployment of simulation-based robotic systems.

### v2

@@ I Introduction
In-context learning has gained traction as a method for adjusting model behavior without gradient updates, widely used in Natural Language Processing (NLP) [16] and recently applied in robotics to improve generalization [17, 18, 19]. For example, Xu et al. [19] enhanced the Decision Transformer (DT) [20] by using new task demonstrations as prompts for online adaptation. Most current in-context learning approaches focus on adapting policies when rewards or expert demonstrations changes, assuming fixed environment dynamics [18, 21]. While different dynamics could be framed as diverse tasks in a multi-task RL setting, more than it’s a counter-intuitive setting, it also becomes impractical with a high-dimensional continuous environment parameter space, requiring many tasks to capture the full range of behaviors. In this paper, we explore a novel question: “Can we adapt simulation environment parameters using the in-context learning paradigm?” Our goal is to eliminate the optimization loop in SysID, in order to accelerate the parameter estimation process by incorporating the in-context learning ability of transformer models.
We introduce in-Context AdaPTation modUle for sim-to-REal system identification, or CAPTURE, to bridge the sim-to-real gap. CAPTURE aims to dynamically adjust the environment parameters online to align simulated and real-world trajectories using next-token prediction based on past interaction data, which includes simulated trajectories, actions, environment parameters, and real-world trajectories as shown in Figure 1. CAPTURE frames the SysID problem as an in-context learning formulation, treating interaction histories as “context.” Unlike existing techniques [4, 22, 23] that rely on short state-action history, CAPTURE aims to learn the complex SysID search process itself through rich and multi-episodic interaction history data. Beyond learning single-step expert parameter matching behaviors, longer interaction histories enables the learned SysID causal transformer to capture a better dynamics representation of the environments. By incorporating in-context learning, CAPTURE provides a smoother and more accurate prediction of subsequent environmental parameters and dynamic behaviors.
@@ III Methodology
Rather than directly adapting the task policy, we prioritize leveraging historical data—including past environment parameters, task state trajectories, and task actions—to estimate next-iteration environment parameters. This approach aims to align simulation dynamics with real-world performance. We assume that as the discrepancy between simulation and real-world environment parameters decreases, the sim-to-real performance gap will naturally narrow. This process is guided by the underlying monotonic properties of the environment parameter adjustments. We start with a description of the problem formulation in Section III-A. Following with three key modules in our pipeline: Section III-B describes the task policy training, Section III-C describes how we generate efficient source-to-target adaptation iterations, and Section III-D defines different components in the SysID causal transformer structure. The main components of CAPTURE pipeline is demonstrated in Figure 2, where it consists the data generation, model training, and inference pipeline.
@@ III-A Problem Formulation
In this section, we define the objective of our approach: aligning simulation and real-world dynamics performance through predicting accurate real-world environment parameters. We begin by introducing the simulation parameters, followed by the task policy and data generation notations, and conclude with the SysID causal transformer notations for domain adaptation.
Environment Parameter Space. We define the task-related environment parameter space $\epsilon\in\mathcal{E}$, which encapsulates different environment parameter values such as the center of mass and sliding friction. We also assume that the tunable environment parameter space $\mathcal{E}$ in simulation is finite and bounded, encompassing properties of different objects. We modify the environment parameters with Robosuite [42], which provides API for modifying the environment parameters through Python code.
SysID Causal Transformer and Interaction Histories. During the SysID causal transformer and data collection, we treat previous SysID iterations as context, including simulated state trajectories $\tau^{sim}=\{s^{sim}_{0},s^{sim}_{1},\dots,s^{sim}_{T}\}$, real state trajectories $\tau^{real}=\{s^{real}_{0},s^{real}_{1},\dots,s^{real}_{T}\}$, rollout action $a\sim\pi(a|s_{0},\epsilon)$, and the past environment parameters $\epsilon$. A robust SysID process must be capable of exploring complex parameter behaviors. To achieve this, we leverage a causal transformer designed to infer environment parameters based on historical interaction data. Following [17], we treat these sequential interactions as history, where current environment parameters depend on previous SysID iterations. Formally, we define the history as:
@@
where $h_{i}$ is the history containing the past $k$ iterations at $i$-th iteration. Our goal is to learn a causal transformer such that it can replicate the SysID process given history. We define the SysID causal transformer, $P_{\theta}$, with the objective of modeling the distribution of simulation parameters conditioned on the history. This approach aims to distill the SysID algorithm through interaction histories, such that the simulated trajectories $\tau^{sim}$ aligns with the real-world trajectories $\tau^{real}$, thereby bridging the sim-to-real gap. The goal of distilling SysID algorithm is to learn the underlying search capabilities from this process by predicting the next iteration in the history. The optimization objective can be formalized as:
@@ III-B Environment-Conditioned RL Training
The environment-conditioned RL task policy $\pi(a\mid s_{0},\epsilon)$ is trained to adapt to varying environment parameters $\epsilon\in\mathcal{E}$. For each episode, $\epsilon$ is sampled uniformly from the parameter space $\mathcal{E}$. Within the episode, the agent selects an action $a$ from $\pi(a\mid s_{0},\epsilon)$, considering the initial state $s_{0}$ and current simulation environment parameter $\epsilon$. This action is executed, producing a state trajectories $\{s_{1},s_{2},\ldots,s_{T}\}$ and a reward $r$. Each episode $\{a,r,s_{0},\epsilon\}$, is stored in the replay buffer. After certain episodes, the policy is updated using Soft Actor-Critic (SAC) [43], refining actions for smoother domain adaptation with predicted parameters.
@@ III-C Source-to-Target SysID Iteration Generation
In the data generation process, we developed source-to-target adaptation transitions that mimic sim-to-real adaptation. Each iteration includes four elements: the current simulation environment parameter $\epsilon_{i}$, the rollout action $a_{i}$, the simulated trajectories $\tau^{source}_{i}$, and the collected target environment trajectories $\tau^{target}_{i}$ under the same action $a_{i}$. The trajectories and actions are obtained through simulation rollouts using an environment-conditioned task policy. Although random action can be used here to collect training data, task policy is preferred to ensure sufficient task-related environment interactions.
In simulation, both source and target values are known, allowing for direct single-step mapping from source to target. However, this approach often performs poorly in real-world deployment when the target’s dynamics representation (state trajectories) lacks sufficient detail. Rather than learning a single-step mapping, we focus on learning a search algorithm that finds the target environment parameter with dynamic representations. The duration of the parameter iteration history $L$ indicates the number of iterations that we pre-defined to generate a complete transition sequence from $\epsilon^{source}$ to $\epsilon^{target}$. We pick a transition number $L=7$ during data generation.
@@
2:Choose parameter transition iteration length $L$
@@ III-D SysID Causal Transformer
To model the sequential relations. We adopted the architectural structure from DT[20], which uses GPT-2[45] as our transformer backbone. Given the collected SysID parameter transition histories, $\mathcal{D}$, our goal is to distill the binary search process through parameter transition sequences with length $L$, where each iteration represents an adaptation iteration. The model predicts the next environment parameter $\hat{\epsilon}_{i+1}$ at iteration $i$ using a next-token prediction framework with a shifted input setup [45]. We sample a multi-episode window of size $k$ from $\mathcal{D}$, where $k$ is a subsequence of the full $L$ iterations. The SysID causal transformer processes this history to predict the next-iteration environment parameter. Each iteration block contains $2+2T$ tokens: one action, one parameter, and $T$ state trajectory tokens for both simulated and real rollouts.
@@ IV-A Experimental Setups
In table air hockey, we test CAPTURE with a higher-dimensional parameter and action space [15, 48]. This task requires precise control and adaptability to match simulated and real-world dynamics. Tunable environment parameters are listed in Table I, with setups shown in Figure 4.
@@ IV-B Baselines and Ablations
Ablations on Different Data Generation Approaches. We modify the data generation module to demonstrate the effectiveness of our distilled searching algorithm over others, including linear interpolation (linterp) and the standard binary search method without randomness (binary), while selecting the next iteration parameters. Linear interpolation randomly sample $L$ points between source and target environment parameters and orderly constructs the SysID transition. The standard binary search method [49] follows a similar setting as ours. However, it does not consider the random beta distribution, it only selects the middle point between the upper and lower bound.
@@ IV-C Sim-to-Sim SysID Evaluation
Object Scooping Sim-to-Sim Evaluation. In the sim-to-sim transfer, we evaluated the normalized context differences, which are one-dimensional in this setting, as shown in Figure 5. Since we use an angle classifier to reduce dependence on accurate sensor data during real-world deployment, reporting trajectory differences becomes irrelevant, as the trajectory in this context is represented solely by a label. Instead, we measure the task’s success rate, defined as lift the object with label (0). Figure 5 shows that CAPTURE achieves a success rate 50% higher than other SysID methods and 70% higher than the DR approach. This is expected, as the baselines lack historical interaction data, making identification only dependent on current scooping points. In contrast, CAPTURE uses a rich previous interaction history, allowing it to gradually narrow down the center of mass search space.
@@
Table Air Hockey Sim-to-Sim Evaluation. In Figure 6, CAPTURE offers better parameter estimation with more stable and accurate adaptation curves. In scenarios where environment parameters require rollout histories, baselines struggle due to their inability to account for historical interactions. For instance, while the ED method might successfully detect the left wall after hitting it, it tends to forget earlier right wall interactions. This short-term memory leads to faster adaptation in simple environments but falls short in more complex ones. In dynamic settings, where SysID needs to identify parameters on both sides for sustained task performance, maintaining a history of parameter updates becomes critical, as it informs subsequent iterations.
@@
Ablation results. In object scooping experiments, we observe that the linear interpolation approach converges more slowly in terms of adaptation iterations, as shown in Figure 5. Due to limited exploration, it hinders performance. Figure 6 shows that CAPTURE + linear interpolation follows a near-linear sim-to-real transition until iteration 7, closely matching the dataset’s transition history. However, it struggles to establish a robust search process due to overfitting to linear interpolated transition histories. Except for left damping parameter, no significant performance differences are seen between randomized binary search (ours) and standard binary search. With added randomized, it did not hinder the estimation performance, whereas it learned a more robust adaptation process.
@@ V Conclusion
This paper introduces a novel in-context learning approach to bridge the sim-to-real gap in robotic tasks by adjusting environment parameters online. By leveraging interaction histories as context, we enable dynamics adaptation to real-world environments without requiring model updates. Evaluated in scooping and table air-hockey tasks, our method outperforms traditional approaches such as domain randomization and TuneNet, reducing the sim-to-real gap and improving both sim-to-sim and sim-to-real performance. The approach leverages historical multi-episode data to infer system parameters and provide a better real-world dynamics prediction. While our method demonstrates strong performance, further alignment is required between state trajectories and environment parameter spaces to extend the approach to multi-task settings, enabling one SysID model to handle multiple task environments. Nonetheless, the framework provides a more efficient and accurate solution for the real-world deployment of simulation-based robotic systems.